OpenStudy (anonymous):

A rectangular pasture has a fence around the perimeter. The length of the fence is 16x7 and the width is 48x4. What is the area of the pasture? (1 point) a) 3x^3 b)128x^11 c)768x^11 d)768x^28 I KNOW the answer is c, but i need it explained. I already found it here once but can ANYONE explain how to do this?

hartnn (hartnn):

Area of rectangular (here pasture) = Length * Width

OpenStudy (anonymous):

Allright, so how do i do it? 16 * ^7 gives nothing and im so confused How do i do 16x^7 * 48x^4? I dont get this problem

OpenStudy (anonymous):

Same thing with this > Simplify. 5^–1(3^–2) If i type this into a calculator i get a stupid number

hartnn (hartnn):

ok, treat constants and variables separately. \(16x^7 \times 48x^4 = (16\times 48) \times(x^7 \times x^4)\) got this ?

OpenStudy (anonymous):

Okay... yes i think i get that... but where did the two Z come from

hartnn (hartnn):

Z ?? there are no Z's there....

OpenStudy (anonymous):

Whoops, sorry, okay, go on. I understand now

OpenStudy (anonymous):

I think i get it 16 * 48 = 768 And 7 + 4 = 11?

hartnn (hartnn):

good!

OpenStudy (anonymous):

Okay, 1 more?

hartnn (hartnn):

actually, its \(x^7 \times x^4 = x^{7+4}=x^{11}\) and sure :)

OpenStudy (anonymous):

Okay something like this Simplify (4xy2)3(xy)5 and (sorry) this Evaluate a–4b2 for a = –2 and b = 4

OpenStudy (anonymous):

Simplify (4xy^2)^3(xy)^5 Sorry

hartnn (hartnn):

i believe the first one looks like this : \((4xy^2)^3(xy)^5\)

hartnn (hartnn):

ok, so you need to remember/understand few thing before u start: \(\huge (ab)^n = a^nb^n\) and \(\huge (a^m)^n=a^{mn}\)

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

The other problem i can't figure out how to write an equation

hartnn (hartnn):

\(\large Evaluate \: \: a^{–4}b^2 \:\: for\:\: a = –2 \:\:and\:\: b = 4\) right ?

OpenStudy (anonymous):

Here you go

hartnn (hartnn):

ok, for the previous one : using, \(\huge (ab)^n = a^nb^n\) \(\large (4xy^2)^3= 4^3 x^3 (y^2)^3\) got this ?

OpenStudy (anonymous):

Oh man. So confused

OpenStudy (anonymous):

Are you doing Simplify (4xy^2)^3(xy)^5 ?

hartnn (hartnn):

yes.

OpenStudy (anonymous):

And you changed that to (4xy^2)^3=4^3x^3(y^2)^3....

OpenStudy (anonymous):

How did you do that... that hurts my brain

hartnn (hartnn):